# PCS Workflow for Veridical Data Science in the Age of AI

### v1

@@ 6 DSLC Stage #4: Predictive Modeling
In the context of statistical modeling, it’s important to draw a distinction between two core concepts: predictions and inference. Predictions refer to the task of predicting responses of unobserved observations based on a trained model, while inference involves the estimation of unknown population parameters. The key difference is that (parameter) inference typically relies on distributional assumptions that may not checkable. The development of PCS thus far has focused on predictions rather than inference. Within the domain of predictions, the PCS framework can be divided into two distinct categories: model training and generating predicted responses. The first category focuses on ensuring the model itself is aligned with reality and stable throughout its development, while the second revolves around using the trained model to attain reliable predictions. For more details on specific types of predictive models, we refer the readers to Chapters 8-12 of Yu and Barter [6].
@@ 6.1 PCS Training of Models
We argue that similar to selecting a measure for accuracy, data scientists should choose a stability metric beforehand which aligns with the specific prediction task. Stability metrics allow practitioners to assess how consistent the model’s performance is when subjected to different data perturbations. Data perturbations are a powerful tool for ensuring stability in model selection. By applying perturbations to the input data (such as adding noise or making small adjustments), one can test how sensitive the model is to changes in the data. For more information on how to perform perturbations to evaluate model stability, we refer the reader to Yu and Kumbier [5].
@@ 6.2 PCS Prediction Results
After completing model training, we have a collection of several “predictive fits”. These fitted algorithms are the result of different judgment calls, including distinct training sets, hyperparameter choices, and more. The data scientist now must choose how predictions will be generated. If the applied problem requires very clear and transparent predictions, then selecting a single model from this collection would be the most appropriate choice. However, there have been recent calls to abandon the oft-used single model framework in favor of ensembles which may pick up on different types of signal [33]. To form such an ensemble, a handful of predictive fits should be selected via a prediction-check on held-out data. A suitable threshold for “good” performance may change based on the application, but possible rules could be to select the top $k\%$ of models or to select all models with some small $\varepsilon$ accuracy of the best-performing model. This type of ensemble is aligned with Leo Breiman’s idea of the Rashomon Effect, which describes the phenomenon where many equally-performing models exist for the same task [13].
@@
Given either a single predictive fit or a “PCS ensemble” consisting of well-performing models, we now must generate predictions for new data points. Obtaining a point estimate is typically straightforward. However, a single predicted value does not display the entire picture. We argue that when possible, some form of uncertainty quantification should be provided alongside the point estimate. In particular, we recommend using PCS prediction pertubation intervals (PPIs) [6, 17]. PPIs represent a range of plausible predictions from models which passed the predictability-check. For single-model use cases, this can correspond to models trained on different training sets, with possible perturbations including bootstrapped samples or different upstream judgment calls. Prediction tasks with a continuous outcome can then take a lower and upper quantile (e.g. 0.05 and 0.95) from the predicted responses to create the PPI. For more information on PPIs, including calibration and coverage information, we refer the reader to Chapter 13 of Yu and Barter [6] and Agarwal et al. [17].
@@
Example. The ciTBI task provides a good example of a case where creating an ensemble is not feasible for the applied problem. Recall that the goal was to create a CDR to help emergency department clinicians make decisions regarding treatment. These CDRs must be incredibly interpretable and easy-to-use, as these decisions are made quickly by the clinician on duty. For the problem of recommending CT scans to screen for ciTBI, the researchers arrived at an inherently explainable tree, which is visualized in Figure 3 of Kuppermann et al. [25]. Ensembling this tree with other predictive fits would make it harder to apply quickly and explain to patients, making the single-model approach better adapted for this task.
[…]
@@ 7 DSLC Stage #5: Evaluation of Results
In practice, scrutinizing your findings should be practiced throughout the DSLC. However, due to its importance, we give it special consideration here. When evaluating the findings from data analysis or results of predictions, it is important to be wary of confirmation bias. Intentional measures should be taken to ensure confirmation bias is not changing how the results are interpreted. One suggestion is to use “dummy” labels when comparing results to competing methods or approaches. Depending on the type of finding, it may also be useful to show “fake” versions of results to domain experts alongside the actual results, allowing the experts to identify that the real findings make sense [6]. This allows collaborators, domain experts, and the researchers to determine that the actual conclusions are trustworthy and offer a noticeable improvement over competing approaches.
@@
[…]
@@ 8 DSLC Stage #6: Communication of Results
The extra effort necessary to make this happen can be a lot of work for data scientists, especially considering many do not have the skills or experienced needed to make easy-to-use software. We recommend data scientists in large organizations or companies collaborate with staff software engineers to create better tools for the adoption of results and findings. For others without such resources on hand, an attainable goal is to create a well-documented GitHub repository with all relevant code, perhaps paired with an interactive web application using tools such as Shiny [34].
@@
[…]
@@ 9 Case Study: Downstream Effects of Data Cleaning
Uncertainty stemming from different data cleaning processes is underemphasized in data science. To demonstrate how choices in data cleaning can change results in other DSLC stages, we explore a real-world case study  [^3] Inspired by Omer Ronen, who performed a similar study as the TA for the PhD-level applied statistics course at UC Berkeley.. During the Spring 2025 semester, the second author taught a graduate statistics course (STAT 214) at UC Berkeley called “Data Analysis and Machine Learning for Real-World Decision Making”. The first author served as a TA with two others. The main task of the first class project was to thoroughly explore and clean the PECARN ciTBI dataset. All students received the same documentation and instructions, including advice from Dr. Aaron Kornblith, a pediatric emergency medicine doctor affiliated with PECARN. Using their cleaned data, the students were then asked to model a potential decision rule for recommending CT scans. Similar to the goal of Kuppermann et al. [25], these recommendations should continue providing scans for children at risk of ciTBI while minimizing the total number of CT scans given. Of the 57 students enrolled in this course, 19 gave us explicit permission to use their cleaned datasets in this work.
@@ 9.2 Judgment Calls
Therefore, our primary focus will be the treatment of missing and invalid data, encapsulating dropping or imputing observations as well as selecting particular covariates. We observe in Figure 2 that the students took wildly different approaches when handling these cases, with some students selecting less than 20% of the features to retain and others dropping nearly a third of the patients from the dataset.
@@ 9.2.1 The Treatment of Missing Values Contributes to Predictive Instability
To evaluate the ramifications of data dropping and imputation, we implement the clinical decision rule (CDR) proposed by Kuppermann et al. [25]. For children under two years of age, this CDR recommends a CT scan if there are signs of altered mental status or basiliar skull fracture. For children two and older, CT is recommended when there are signs of altered mental status or palpable skull fracture. This decision rule is selected to evaluate results because it uses a small set of variables which all 19 students kept.
@@ 9.2.2 The Selection of Covariates for Modeling Increases Predictive Uncertainty
Figure 2 shows that students also made a wide variety of choices when deciding which variables may be helpful for downstream modeling. Since the CDR only uses a small number of variables which each cleaned dataset has in common, it does not capture the uncertainty resulting from the inclusion of different features. To investigate this uncertainty, we fit a logistic regression model to each student’s cleaned dataset. Specifically, we use a 70/30 train-test split, fitting the regression model to the training set and evaluating performance on the held-out test set. We set the threshold for CT scan recommendation at 1/313. This number is selected since it represents the lifetime attributable risk of cancer due to head CT scans on children [24]. Therefore, this threshold corresponds to recommending CT if the probability of ciTBI is greater than the lifetime attributable risk of cancer, and avoiding CT otherwise. We note that because logistic regression cannot be performed in the presence of missingness, four cleaned datasets had to be further preprocessed. For these datasets, we first drop all columns with more than 10,000 NA values. These dropped features largely represent descriptors for other covariates. We then drop any remaining patients with incomplete entries.
@@ 10 Discussion
The PCS workflow for veridical data science (VDS) provides a practical approach to create replicable results through systematically addressing uncertainty caused by judgment calls in the DSLC. By incorporating tools and practices that check reality and quantify both stability and uncertainty in a computationally efficient manner, PCS enables more reliable and truthful data science. The PCS principles are broadly applicable across domains where data science and AI are used to draw conclusions, such as the social sciences and biomedical sciences. Although this paper focuses on synthesizing the current practice of PCS while adding GenAI guidance, the framework has many possible extensions. Two such extensions are discussed briefly below.
@@
PCS Design. One active research direction involves creating PCS-adherent strategies for generalized experimental design [35]. Whereas traditional experimental design often relies on expensive and time-consuming controlled experiments to establish causality, PCS design emphasizes carefully structuring the DSLC to arrive at cost-effective decisions. For instance, Wang et al. [11] is able to identify genetic drivers of cardiac hypertrophy by providing interpretable hypothesis generation recommendations for further experiments. Further works may include treatment effect estimation using prediction-checked adjustments or extending PCS to Bayesian design settings.
@@
PCS Methodology. The PCS framework encourages practitioners to select reality and stability-checked models. While certain out-of-the-box models like random forests may sometimes provide predictive stability, certain applications may require developing new inherently stable and interpretable methodology. Examples include Dwivedi et al. [10], which introduces a method for discovering stable subgroups in randomized experiments, and Basu et al. [18], which proposes an algorithm for identifying predictive and stable high-order interactions. Any applied problem with a large amount of uncertainty from judgment calls may benefit from similar PCS-aligned modeling algorithms.
@@
We hope that this paper will encourage practitioners and researchers to try the PCS framework for their problems of interest. For those who would like to learn more, we recommend watching the recordings of the 2024 Berkeley-Stanford Workshop on VDS or gathering information on other workshops such as the June 2025 Rome Workshop on VDS and July 2025 VDS Workshop for Biology. Materials (including code) for incorporating PCS principles into data-related project can be found on the Yu Group website and GitHub. New PCS developments can also be found at second author’s website.

### v2

@@ 6 DSLC Stage #4: Predictive Modeling
In the context of statistical modeling, it’s important to draw a distinction between two core concepts: predictions and inference. Predictions refer to the task of predicting responses of unobserved observations based on a trained model, while inference involves the estimation of unknown population parameters. The key difference is that (parameter) inference typically relies on distributional assumptions that may not checkable. The development of PCS thus far has focused on predictions rather than inference. Within the domain of predictions, the PCS framework can be divided into two distinct categories: model training and generating predicted responses. The first category focuses on ensuring the model itself is aligned with reality and stable throughout its development, while the second revolves around using the trained model to attain reliable predictions. For more details on specific types of predictive models, we refer the readers to Chapters 8-12 of veridical2024yu.
@@ 6.1 PCS Training of Models
We argue that similar to selecting a measure for accuracy, data scientists should choose a stability metric beforehand which aligns with the specific prediction task. Stability metrics allow practitioners to assess how consistent the model’s performance is when subjected to different data perturbations. Data perturbations are a powerful tool for ensuring stability in model selection. By applying perturbations to the input data (such as adding noise or making small adjustments), one can test how sensitive the model is to changes in the data. For more information on how to perform perturbations to evaluate model stability, we refer the reader to vds2020yu.
@@ 6.2 PCS Prediction Results
After completing model training, we have a collection of several “predictive fits”. These fitted algorithms are the result of different judgment calls, including distinct training sets, hyperparameter choices, and more. The data scientist now must choose how predictions will be generated. If the applied problem requires very clear and transparent predictions, then selecting a single model from this collection would be the most appropriate choice. However, there have been recent calls to abandon the oft-used single model framework in favor of ensembles which may pick up on different types of signal [rudin2024models]. To form such an ensemble, a handful of predictive fits should be selected via a prediction-check on held-out data. A suitable threshold for “good” performance may change based on the application, but possible rules could be to select the top $k\%$ of models or to select all models with some small $\varepsilon$ accuracy of the best-performing model. This type of ensemble is aligned with Leo Breiman’s idea of the Rashomon Effect, which describes the phenomenon where many equally-performing models exist for the same task [breiman2001cultures].
@@
Given either a single predictive fit or a “PCS ensemble” consisting of well-performing models, we now must generate predictions for new data points. Obtaining a point estimate is typically straightforward. However, a single predicted value does not display the entire picture. We argue that when possible, some form of uncertainty quantification should be provided alongside the point estimate. In particular, we recommend using PCS prediction perturbation intervals [veridical2024yu, agarwal2025pcsuq]. Prediction perturbation intervals represent a range of plausible predictions from models which passed the predictability-check. For single-model use cases, this can correspond to models trained on different training sets, with possible perturbations including bootstrapped samples or different upstream judgment calls. Prediction tasks with a continuous outcome can then take a lower and upper quantile (e.g. 0.05 and 0.95) from the predicted responses to create the interval. For more information on prediction perturbation intervals, including calibration and coverage information, we refer the reader to Chapter 13 of veridical2024yu and agarwal2025pcsuq.
@@
Example. The ciTBI task provides a good example of a case where creating an ensemble is not feasible for the applied problem. Recall that the goal was to create a CDR to help emergency department clinicians make decisions regarding treatment. These CDRs must be incredibly interpretable and easy-to-use, as these decisions are made quickly by the clinician on duty. For the problem of recommending CT scans to screen for ciTBI, the researchers arrived at an inherently explainable tree, which is visualized in Figure 3 of kuppermann2009tbi. Ensembling this tree with other predictive fits would make it harder to apply quickly and explain to patients, making the single-model approach better adapted for this task.
ȷ  Generative AI Tips & Warnings  . Do not allow GenAI to create the entire modeling pipeline at once. Instead, break it into smaller modules, such as parameter tuning and model selection. Always review the generated code for correctness, efficiency, and alignment with the problem context. Tools may forget to use random seeds, harming reproducibility of predictions.
Key Takeaways & Action Items Stability is encouraged in model development through perturbations to both data and algorithms. Rather than relying on a single “optimal” model, the PCS framework encourages using multiple models when feasible. The well-performing models can be combined into a “PCS ensemble” to create more stable point estimates as well as prediction perturbation intervals.
@@ 7 DSLC Stage #5: Evaluation of Results
In practice, scrutinizing your findings should be practiced throughout the DSLC. However, due to its importance, we give it special consideration here. When evaluating the findings from data analysis or results of predictions, it is important to be wary of confirmation bias. Intentional measures should be taken to ensure confirmation bias is not changing how the results are interpreted. One suggestion is to use “dummy” labels when comparing results to competing methods or approaches. Depending on the type of finding, it may also be useful to show “fake” versions of results to domain experts alongside the actual results, allowing the experts to identify that the real findings make sense [veridical2024yu]. This allows collaborators, domain experts, and the researchers to determine that the actual conclusions are trustworthy and offer a noticeable improvement over competing approaches.
@@
ȷ  Generative AI Tips & Warnings  . GenAI can help create dummy labels and altered plots for comparison. Coding assistants can generate different types of plots to verify consistency. Do not allow LLMs to draw conclusions from any visualizations. This is the job of the analyst and collaborating domain experts.
Key Takeaways & Action Items Compare results to competing methods and approaches without knowing which is which, thus helping avoid confirmation bias. Ensure that results are stable across reasonable visualization schemes.
@@ 8 DSLC Stage #6: Communication of Results
The extra effort necessary to make this happen can be a lot of work for data scientists, especially considering many do not have the skills or experienced needed to make easy-to-use software. We recommend data scientists in large organizations or companies collaborate with staff software engineers to create better tools for the adoption of results and findings. For others without such resources on hand, an attainable goal is to create a well-documented GitHub repository with all relevant code, perhaps paired with an interactive web application using tools such as Shiny [chang2024shiny].
@@
ȷ  Generative AI Tips & Warnings  . Language models can draft documentation for code and/or data repositories. Individuals with limited programming experience can use coding assistants to create simple web applications with demonstrations of their conclusions. As always, GenAI produced documentation must be checked and verified.
Key Takeaways & Action Items Findings should be presented in a way that makes them easy to understand and apply—whether it’s through presentations, software, or publications.
@@ 9 Case Study: Downstream Effects of Data Cleaning
Uncertainty stemming from different data cleaning processes is underemphasized in data science. To demonstrate how choices in data cleaning can change results in other DSLC stages, we explore a real-world case study  [^3] Inspired by Omer Ronen, who performed a similar study as the TA for the PhD applied statistics course at UC Berkeley., [^4] All code and data can be found at https://github.com/Yu-Group/pcs-workflow.. During the Spring 2025 semester, the second author taught a graduate statistics course (STAT 214) at UC Berkeley called “Data Analysis and Machine Learning for Real-World Decision Making”. The first author served as a TA with two others. The main task of the first class project was to thoroughly explore and clean the PECARN ciTBI dataset. All students received the same documentation and instructions, including advice from Dr. Aaron Kornblith, a pediatric emergency medicine doctor affiliated with PECARN. Using their cleaned data, the students were then asked to model a potential decision rule for recommending CT scans. Similar to the goal of kuppermann2009tbi, these recommendations should continue providing scans for children at risk of ciTBI while minimizing the total number of CT scans given. Of the 57 students enrolled in this course, 19 gave us explicit permission to use their cleaned datasets in this work.
@@ 9.2 Judgment Calls
We observe in Figure 2 that the students took wildly different approaches when handling these cases, with some students selecting less than 20% of the features to retain and others dropping nearly a third of the patients from the dataset. Therefore, our primary focus will be the treatment of missing and invalid data, encapsulating dropping or imputing observations as well as selecting particular covariates. While ensembling these heterogeneously processed datasets would provide an avenue for improved predictions, the varying approaches to missingness creates challenges for combined inference and validation. Chapter 13 of veridical2024yu details the construction of PCS ensembles and prediction perturbation intervals for differing workflows, but more work is needed to properly combine estimators fit on similar but distinct datasets.
@@ 9.2.1 The Treatment of Missing Values Contributes to Predictive Instability
To evaluate the ramifications of data dropping and imputation, we implement the clinical decision rule (CDR) proposed by kuppermann2009tbi. For children under two years of age, this CDR recommends a CT scan if there are signs of altered mental status or basiliar skull fracture. For children two and older, CT is recommended when there are signs of altered mental status or palpable skull fracture. This decision rule is selected to evaluate results because it uses a small set of variables which all 19 students kept.
@@ 9.2.2 The Selection of Covariates for Modeling Increases Predictive Uncertainty
Figure 2 shows that students also made a wide variety of choices when deciding which variables may be helpful for downstream modeling. Since the CDR only uses a small number of variables which each cleaned dataset has in common, it does not capture the uncertainty resulting from the inclusion of different features. To investigate this uncertainty, we fit a logistic regression model to each student’s cleaned dataset. Specifically, we use a 70/30 train-test split, fitting the regression model to the training set and evaluating performance on the held-out test set. We set the threshold for CT scan recommendation at 1/313. This number is selected since it represents the lifetime attributable risk of cancer due to head CT scans on children [smith-bindman2025cancer]. Therefore, this threshold corresponds to recommending CT if the probability of ciTBI is greater than the lifetime attributable risk of cancer, and avoiding CT otherwise. We note that because logistic regression cannot be performed in the presence of missingness, four cleaned datasets had to be further preprocessed. For these datasets, we first drop all columns with more than 10,000 NA values. These dropped features largely represent descriptors for other covariates. We then drop any remaining patients with incomplete entries.
@@ 10 Discussion
The PCS workflow for veridical data science (VDS) provides a practical approach to create replicable results through systematically addressing uncertainty caused by judgment calls in the DSLC. By incorporating tools and practices that check reality and quantify both stability and uncertainty in a computationally efficient manner, PCS enables more reliable and truthful data science. The PCS principles are broadly applicable across domains where data science and AI are used to draw conclusions, such as the social sciences and biomedical sciences. Although this paper focuses on synthesizing the current practice of PCS while adding GenAI guidance, the framework is still evolving and has many possible extensions. Four such extensions are discussed briefly below.
(1) Mitigating Difficulties of Implementing the PCS Workflow. While we maintain that the PCS principles are universally applicable, it is important to acknowledge that implementing the current workflow can require a non-trivial amount human effort and computational resources. Although some implementations can be rather straightforward, such as obtaining PCS rankings of gene expression features [wang2024epistasis], keeping track of results under different judgment calls grows exponentially [gelman2013garden]. For this reason, adhering to the entire PCS workflow may not be feasible in certain settings. Specialized software which tracks and manages judgment calls would make the implementation of PCS much easier. The Yu Group has begun addressing this bottleneck [duncan2022veridicalflow, Duncan2024simChef, elliott2024designing]. We are optimistic about a wider adoption of PCS in the near future while recognizing that more work is needed to fully overcome these challenges.
(2) PCS GenAI Benchmarking. As GenAI capabilities continue to advance, companies and researchers are beginning to develop data science ‘agents’ that can automate complex data analysis tasks and even generate actionable conclusions [sun2025agenticdataagenticdataanalytics, fine2025datascienceagent]. This has resulted in a growing body of benchmarking research which aims to evaluate the performance of these agents on real-world data science workflows [gu2024blade, majumderdiscoverybench]. We believe that the PCS workflow naturally lends itself to such a purpose, as agents can be evaluated on the predictability and stability of their findings under different judgment calls as well as the produced code’s adherence to the MERITS criteria [elliott2024designing]. For example, benchmarking could include applying agents to datasets with predefined perturbations or alternate modeling choices, with each agent being scored on the consistency of their conclusions across variations. Furthermore, much effort has been put into developing GenAI tools with ‘mathematical reasoning’ capabilities, but very little effort has been given to statistical reasoning. We see a PCS benchmarking process as having the potential to ensure trustworthy agent findings, as well as a way to evaluate the statistical reasoning abilities of GenAI systems.
@@
(3) PCS Design. One active research direction involves creating PCS-adherent strategies for generalized experimental design [murphy2024modernizing]. Whereas traditional experimental design often relies on expensive and time-consuming controlled experiments to establish causality, PCS design emphasizes carefully structuring the DSLC to arrive at cost-effective decisions. For instance, wang2024epistasis is able to identify genetic drivers of cardiac hypertrophy by providing interpretable hypothesis generation recommendations for further experiments. Further works may include treatment effect estimation using prediction-checked adjustments or extending PCS to Bayesian design settings.
@@
(4) PCS Methodology. The PCS framework encourages practitioners to select reality and stability-checked models. While certain out-of-the-box models like random forests may sometimes provide predictive stability, certain applications may require developing new inherently stable and interpretable methodology. Examples include dwivedi2020stable, which introduces a method for discovering stable subgroups in randomized experiments, and basu2018iterative, which proposes an algorithm for identifying predictive and stable high-order interactions. Any applied problem with a large amount of uncertainty from judgment calls may benefit from similar PCS-aligned modeling algorithms.
@@
We hope that this paper will encourage practitioners and researchers to try the PCS framework for their problems of interest. For those who would like to learn more, we recommend watching the recordings of the 2024 Berkeley-Stanford Workshop on VDS or gathering information on other workshops such as the June 2025 Rome Workshop on VDS and July 2025 VDS Workshop for Biology by QB3 at UC Berkeley. Materials (including code) for incorporating PCS principles into data-related project can be found on the Yu Group website and GitHub. New PCS developments, including a biweekly talk series on Veridical Data Science in Biology (in the planning phase), will be available at the second author’s website.